# Improving the Efficiency of Self-Supervised Adversarial Training through Latent Clustering-Based Selection

### v1

@@ 4.3. Latent Clustering-Based Selection
where $\pi_{j}$ is the mixing coefficient for the $j$-th Gaussian component, and $\mathcal{N}(\bm{z}\mid\bm{\mu}_{j},\mathbf{\Sigma}_{j})$ represents the Gaussian distribution with mean $\bm{\mu}_{j}$ and covariance $\Sigma_{j}$. We denote $\bm{p}(\bm{z})=[p_{1}(\bm{z}),p_{2}(\bm{z}),\ldots,p_{k}(\bm{z})]$ as the probability vector. To identify data points near the decision boundary, we focus on those points for which the posterior probabilities $\bm{p}(\bm{z})$ are similar across multiple Gaussian components, indicating that they are near the boundary between different clusters. Specifically, for each data point $\bm{z}$, we calculate the difference between the highest and the second-highest posterior probabilities: $\Delta p=|p_{1}(\bm{z})-p_{2}(\bm{z})|$, where $p_{1}(\bm{z})$ and $p_{2}(\bm{z})$ are the highest and second-highest posterior probabilities, respectively. Data points with the smallest $\Delta\gamma$ values are selected, as they are more likely to be near decision boundaries. Finally, the top $\alpha|\mathcal{S}_{u}|$ unlabeled data with the smallest $\Delta p$ values are used to form the selected subset. As will be illustrated in Section 5, the robustness performance of SSAT can be largely maintained when using LCS-KM and LCS-GMM to select a small subset of unlabeled data (e.g., $\alpha=10\%$ or $\alpha=20\%$). More detailed discussions of the difference between these data selection schemes are provided in Section 6.1, where we visualize the selected unlabeled data in a two-dimensional latent space.
5. Experiments
In this section, we comprehensively evaluate the performance of our proposed data selection schemes, as described in Algorithms 1 and 2, using two widely recognized image benchmarks: SVHN (Netzer et al., 2011) and CIFAR-10 (Alex, 2009) (Section 5.2). Following existing literature, we consider $\ell_{\infty}$-norm bounded perturbations with $\epsilon=0.015$ on SVHN and $\epsilon=0.031$ on CIFAR-10. We also highlight the computational advantages of our methods, particularly their ability to reduce the convergence time of SSAT algorithms (Section 5.3). Furthermore, we assess the generalizability of our techniques by applying them to a real-world medical dataset, demonstrating the applicability of our methods beyond standard image benchmarks (Section 5.4).
Table 1. Comparison results of self-supervised adversarial training performance under different unlabeled data selection schemes and varying ratios on two image benchmarks. For SVHN, we consider $\ell_{\infty}$ perturbations with $\epsilon=0.015$ and train models on Wide ResNet 16-8. For CIFAR-10, we consider $\ell_{\infty}$ perturbations with $\epsilon=0.031$ and train models on Wide ResNet 28-10.
@@ 5. Experiments
LCS-KM $88.7$ $60.7$ $57.8$ $14.78$ $85.5$ $60.2$ $57.2$ $15.27$
5.1. Experimental Settings
First, we introduce the necessary details to understand our experimental results. All the remaining details for reproducing our experiments are provided in Appendix B.
Dataset. For the image benchmark experiments, the proposed selection schemes are initially applied to external unlabeled datasets, following the implementation protocols outlined by Carmon et al. (2019). For the SVHN experiments, models are trained on $73$K labeled digit images from the original SVHN dataset, supplemented by $531$K additional unlabeled SVHN images. Specifically, the CIFAR-10 dataset comprises $50$K labeled training images and $10$K labeled test images, with models trained using SSAT augmented by an external unlabeled dataset $\mathcal{S}_{u}$ containing $500$K images sampled from the $80$M Tiny Images (80M-TI) dataset. To assess the generalizability of our selection schemes, we further evaluate SSAT algorithms using $1$M synthetically generated images created by the Denoising Diffusion Probabilistic Model (DDPM) for both CIFAR-10 and SVHN, following the protocols described by Gowal et al. (2021).
Configuration. We primarily conduct experiments on SSAT using WideResNet architectures. We train models using TRADES (Zhang et al., 2019) on both labeled and unlabeled data coupled with pseudo labels generated by the intermediate model. In Section 6.3, we evaluate the sensitivity of our methods for varying $\ell_{\infty}$ perturbation size, PGD-based adversarial training with varying attack steps, and $\ell_{2}$ perturbations. To standardize training, an epoch is defined as processing $50$K data points, regardless of the total dataset size. This is achieved by calculating the required number of batches to cover $50$K data points based on the batch size. This approach, adopted in prior works (Carmon et al., 2019; Gowal et al., 2021), ensures that training time per epoch remains consistent across datasets of varying sizes.
SSAT. For experiments with external unlabeled data, SSAT models are trained for a total of $200$ epochs, whereas experiments with generated data extend to $400$ epochs. To evaluate the performance of each data selection scheme, models are saved every $25$ epoch of SSAT, and the model achieving the highest robust accuracy is selected as the “best” model. This approach aligns with the early stopping practices commonly used in adversarial ML literature (Zhang et al., 2019; Rice et al., 2020). In experiments with external data and selection ratios $\alpha\in\{10\%,20\%\}$, we consistently observe peak robust accuracy around $100$ epochs for CIFAR-10 and $75$ epochs for SVHN. Conversely, when no data selection scheme is applied, peak performance is delayed to approximately $200$ epochs. Total training time is reported as the duration required to achieve the best model performance.
Intermediate Model. As a foundational step, the proposed selection schemes require training an intermediate model that utilizes the same architecture as the final model. This intermediate model is trained using standard supervised learning for $100$ epochs, with training on the WRN-28-10 architecture taking approximately $54$ minutes and $43$ seconds. Notably, this time is excluded from the reported training durations presented in the results table. The intermediate model fulfills two critical roles: it not only facilitates the pseudo-labeling of unlabeled data but also constitutes a necessary component for the implementation of any data selection strategy. As such, this training step is integral to the overall workflow, irrespective of whether a specific selection method is employed.
Evaluation Metric. To evaluate model robustness, we primarily focus on using multi-step PGD attacks (PGD) (Madry et al., 2017) due to their wide adoption in existing adversarial ML literature. We also report the robust accuracy using AutoAttack (AA) (Croce and Hein, 2020b) in our main experiments on image benchmarks for more rigorous evaluation and the model’s clean accuracy on normal inputs (Clean) as a reference.
5.2. Main Evaluation on Image Benchmarks
@@
SVHN. Our SVHN evaluation results, summarized in Table 1, demonstrate the efficacy of SSAT algorithms with unlabeled data using various selection schemes and $\alpha\in\{0\%,1\%,10\%,20\%,100\%\}$ trained on a WRN-16-8 architecture. Specifically, employing the LCS-KM selection strategy with $10\%$ or $20\%$ of the additional data achieves robust accuracies of $86.3\%$ and $86.6\%$, respectively, which closely approximate the $86.0\%$ robustness achieved using the entire unlabeled dataset. In contrast, random selection with $20\%$ of the data yields a robust accuracy of only $82.3\%$, underscoring the superior efficacy of informed selection strategies such as LCS-KM. The robust accuracies evaluated using AutoAttack (i.e., the “AA” column of Table 1) demonstrate similar patterns compared to that using multi-step PGD Attacks (i.e., the “PGD” column of Table 1). Additionally, alternative methods, including PCS and LCS-GMM, consistently outperform random selection, reinforcing the robustness and adaptability of the proposed data selection framework. Notably, the robustness of models trained using SSAT peaks when the boundary ratio parameter $\beta$ is inversely proportional to the selection ratio $\alpha$, suggesting that smaller $\beta$ values are favorable as the data inclusion proportion increases. We provide detailed analyses and discussions of these ratio hyperparameter configurations in Section 6.2.
@@
We also note that the use of ground-truth labels reveals negligible differences compared to pseudo-labeled data, indicating that the primary performance improvements stem from the proposed data selection strategy rather than label accuracy (see Table 4 in Appendix D.2 for details). Additionally, experiments with $1$M data points generated using DDPM for SSAT (Gowal et al., 2021) highlight that selecting only $10\%$$20\%$ of data through LCS-KM achieves comparable performance to utilizing the full unlabeled dataset, with LCS-KM consistently demonstrating superior results. Importantly, training SSAT with selected data requires only $75$ epochs to optimize robust accuracy, whereas training with the full dataset demands up to $400$ epochs. This significant reduction in computational cost further underscores the practical advantages of the proposed selection methodology, which will be elaborated in Section 5.3.
@@
CIFAR-10. For CIFAR-10 experiments, we adopt the WRN-28-10 architecture to train models using self-supervised adversarial training algorithms. To assess the generalizability of our observations, we also conduct ablation studies on a ResNet-18 architecture, which are detailed in Table 5 in Appendix D.4). The impact of varying proportions, $\alpha$, of pseudo-labeled unlabeled data was analyzed, revealing that selecting $20\%$ of the external data using LCS-KM achieved a robust accuracy of $60.7\%$, comparable to the $61.2\%$ obtained with the entire dataset. In contrast, the random selection of the same proportion yielded only $57.8\%$, underscoring the advantages of strategic data selection techniques such as LCS-KM. Building on the SVHN experiments, we extend our analysis to CIFAR-10 with synthetically generated data. Using $1$ million images generated by DDPM (Gowal et al., 2021), we discover that utilizing just $20\%$ of the generated data via LCS-KM resulted in performance metrics nearly identical to those achieved with the full dataset. These trends are consistent across experiments with $\ell_{2}$ perturbations (see Figure LABEL:fig:two_graphs_ab for evidence).
@@
Our findings emphasize the robustness and adaptability of the proposed data selection schemes, especially for LCS-KM, even under varying experimental configurations. In terms of computational efficiency, CIFAR-10 experiments highlight the significant advantages of data selection. When leveraging selected external unlabeled data, optimal robust performance was achieved after $100$ epochs of training. By contrast, using the full unlabeled or generated datasets without selection required extended training durations of $200$ and $400$ epochs, respectively, to reach comparable performance. These results validate the efficacy of our latent clustering-based selection approach in enhancing self-supervised adversarial training (SSAT). The proposed method not only maintains the robust accuracy of the final model, $\theta_{\mathrm{final}}$, but also substantially reduces computational costs and accelerates the training process, underscoring its practical utility and scalability in adversarial training frameworks. Table 3 in Appendix C summarizes the comparisons, showcasing that LCS-KM improves model robustness while significantly improving the efficiency of SSAT algorithms with full unlabeled data.
5.3. Computational Benefits

### v2

@@ III-B Strategic Selection via Latent Clustering
where $\pi_{j}$ is the mixing coefficient for the $j$-th Gaussian component, and $\mathcal{N}(\bm{z}\mid\bm{\mu}_{j},\mathbf{\Sigma}_{j})$ denotes the Gaussian distribution with mean $\bm{\mu}_{j}$ and covariance $\Sigma_{j}$. We focus on data points with similar top posterior probabilities, suggesting that they are near the boundary between Gaussian distributions. Specifically, for each $\bm{z}$, we calculate the difference between the highest and the second-highest posterior probabilities: $\Delta p=|p_{1}(\bm{z})-p_{2}(\bm{z})|$, where $p_{1}(\bm{z})$ and $p_{2}(\bm{z})$ are the highest and second-highest posterior probabilities, respectively. Data points with the smallest $\Delta\gamma$ values are selected, as they are more likely to be near decision boundaries. Finally, the top $\alpha|\mathcal{S}_{u}|$ unlabeled data with the smallest $\Delta p$ values are used to form the selected subset. As will be illustrated in Section IV, the robustness performance of SSAT can be largely maintained when using LCS-KM and LCS-GMM to select a small subset of unlabeled data (e.g., $\alpha=10\%$ or $\alpha=20\%$). More detailed discussions of the difference between these data selection schemes are provided in Section V-B, where we visualize the selected unlabeled data in a two-dimensional latent space.
Algorithm 3 SSAT with Guided DDPM Fine-Tuning
1:  Input: labeled $\mathcal{S}_{l}$; pre-trained DDPM $\theta_{\mathrm{pre}}$; fine-tuning epochs $S$; trade-off parameter $\lambda$; other parameters $\epsilon$, $\alpha|\mathcal{S}_{u}|$, $\beta$, $\gamma$, $k$, $\eta$
3:  Initialize $\mathcal{G}_{u}$ as empty set
4:  $\mathcal{G}_{u}\leftarrow$ add $(1-\beta)\cdot\alpha|\mathcal{S}_{u}|$ points generated by pre-trained $\theta_{\mathrm{pre}}$
5:  Initialize $\theta$ as $\theta_{\mathrm{pre}}$
6:  for $s=1,2,\ldots,S$ do
7:   Sample batches of size $b$: $\bm{x}_{0}\sim\mathcal{S}_{l}$, $t\sim\mathcal{U}(T)$, $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\mathbf{I})$
8:   $\bm{x}_{t}\leftarrow\sqrt{\bar{\alpha}_{t_{i}}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$
9:   $L_{\mathrm{DDPM}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{x}_{t},t)\|_{2}^{2}$
10:   if using PCG then
11:    $L_{\mathrm{reg}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\ell_{\mathrm{PC}}(g_{\theta}(\bm{x}_{t},t))\|_{2}^{2}$
12:   else if using LCG-KM then
13:    $L_{\mathrm{reg}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\ell_{\mathrm{KM}}(g_{\theta}(\bm{x}_{t},t))\|_{2}^{2}$
14:   else if using LCG-GMM then
15:    $L_{\mathrm{reg}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\ell_{\mathrm{GMM}}(g_{\theta}(\bm{x}_{t},t))\|_{2}^{2}$
16:   end if
17:   $\theta\leftarrow\theta-\eta\nabla_{\theta}\big(L_{\mathrm{DDPM}}+\lambda\cdot L_{\mathrm{reg}}\big)$
18:  end for
19:  $\mathcal{G}_{u}\leftarrow$ add $\beta\cdot\alpha|\mathcal{S}_{u}|$ points generated by fine-tuned $\theta$
20:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{G}_{u},\gamma)$
21:  Output: generated dataset $\mathcal{G}_{u}$, final model $\theta_{\mathrm{final}}$
III-C Guided Diffusion via Latent Clustering
While our selection-based strategy can significantly reduce the unlabeled data set size and lower the overall training time, it has a critical drawback–when SSAT employs a generative approach, all our previous methods require generating the full synthetic unlabeled dataset as a prerequisite before selection, which is inefficient in terms of both generation time and storage requirement. As shown in Table II, generating 1M CIFAR-10-like synthetic images via DDPM requires around $4$ hours on a single A100 GPU. Note that the best current SSAT method documented in the RobustBench leaderboard [17] utilizes a few hundred million synthetically generated images, suggesting the total generation time can be a significant computational overhead. Therefore, a natural question arises: whether pre-generating synthetic data and then applying strategic selection schemes can be further improved. In this section, we propose an alternative generative pipeline, detailed in Algorithm 3, which answers the above question affirmatively. In particular, we fine-tune a pretrained DDPM using novel guidance losses for more efficient unlabeled data generation based on the insights from previous sections. In this way, we directly generate a small, critical set of boundary-adjacent data points, thereby avoiding redundant data generation and further reducing the computation overhead.
DDPM. Diffusion model resembles a milestone in the history of generative modeling, which is the current state-of-the-art for high-quality image generation and synthesis. A denoising diffusion probabilistic model (DDPM) is a deep neural network trained to learn how to step-by-step reverse the forward process of adding Gaussian noise to an image [23]. By gradually removing Gaussian noise through the reverse process, DDPM can thus generate real-looking images from pure noise. Let $q_{0}$ be a real image distribution over $\mathbb{R}^{d}$ and $\mathcal{N}(\bm{0},\mathbf{I})$ be the noise distribution. The forward process defines how to map images $\bm{x}_{0}\sim q_{0}$ to Gaussian noises $\bm{x}_{T}\sim\mathcal{N}(\bm{0},\mathbf{I})$ through conditional sampling: for any $t\in[T]$, let
$q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t};\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}\mathbf{I}),$ (8)
where $\beta_{1},\beta_{2},\ldots,\beta_{T}$ are scheduling parameters. According to the properties of Gaussian distributions, Equation 8 implies that we can directly sample $\bm{x}_{t}$ for any $t\in[T]$ based on:
$q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\sqrt{\bar{\alpha}_{t}}\bm{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}),$ (9)
where $\bar{\alpha}_{t}=\Pi_{s=1}^{t}(1-\beta_{s})$. Based on Equation 9, Ho et al. [23] proposed to train DDPM by minimizing a regression loss:
$L_{\mathrm{DDPM}}(\theta)=\mathbb{E}_{\bm{x}_{0}\sim q_{0},\bm{\epsilon}\sim\mathcal{N}(\bm{0},\mathbf{I}),t\in\mathcal{U}(T)}\big\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{x}_{t},t)\big\|_{2}^{2},$ (10)
where $\bm{x}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$ denotes the noisy image corresponding to $\bm{x}_{0}$ at $t$-th forward step, $\bm{\epsilon}_{\theta}:\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$ stands for a neural net with weights $\theta$ for noise prediction from $\bm{x}_{t}$, and $\mathcal{U}(T)$ denotes the uniform distribution over the set $\{1,2,\ldots,T\}$. Once the model $\bm{\epsilon}_{\theta}$ is trained, we can then generate synthetic images from Gaussian noise via the following reverse (or denoising) process: for $t=T,\ldots,1$,
$\bm{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\big(\bm{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)\big)+\sqrt{\beta_{t}}\bm{\delta},$ (11)
where $\alpha_{t}=1-\beta_{t}$, and $\bm{\delta}\sim\mathcal{N}(\bm{0},\mathbf{I})$. For simplicity, denote by $g_{\theta}(\bm{x}_{t},t)$ the right-hand side of Equation 11, and $g_{\theta}:\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$ can be viewed as one-step DDPM denoising.
Define DDPM Guidance Loss. Now that we have explained DDPM’s forward and backward processes, we can introduce the detailed design of our fine-tuning schemes. Borrowing the insights from previous sections, our method aims to fine-tune a DDPM model to directly generate the required subset of boundary-adjacent unlabeled data points for more efficient SSAT. Similar to our selection-based methods, we consider three designs: prediction confidence-guided generation (PCG), latent clustering-guided generation with k-means (LCG-KM), and latent clustering-guided generation with Gaussian mixture models (LCG-GMM). Specifically, we define a guidance loss $\ell_{\mathrm{guide}}:\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ corresponding to each design, which will serve as a fine-tuning regularization term to bias the diffusion model to generate more boundary-adjacent samples.
PCG. The first design utilizes the prediction confidence of the intermediate model $f_{\hat{\theta}}$ and defines the guidance loss as
[…]
A larger value of $\ell_{\mathrm{PC}}(\bm{x})$ suggests that $\bm{x}$ is more likely to be closer to the model’s decision boundary. As we’ve explained in Section III-B, the prediction confidence of a neural network classifier can be an unreliable estimator. Therefore, we propose alternative guidance losses using more robust latent clustering-based approaches. Based on different clustering techniques, we again introduce two possible variants.
LCG-KM. For LCG-KM, we first obtain the latent embeddings of all the labeled data with respect to the intermediate model, namely $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{l}\}$, and then apply k-means clustering algorithm to partition the embeddings into $k$ clusters. Note that this step is different from LCS-KM, because we now do not pre-generate the unlabeled dataset. Let $\{\bm{\mu}_{1},\bm{\mu}_{2},\ldots,\bm{\mu}_{k}\}$ denote the centroids corresponding to the $k$ returned clusters. We define the guidance loss as:
[…]
where $d_{1}(\bm{x};h_{\hat{\theta}})=\min_{j\in[k]}\|h_{\hat{\theta}}(\bm{x})-\bm{\mu}_{j}\|_{2}$ denotes the Euclidean distance between the latent embeddings of $\bm{x}$ and the closest centroid, $d_{2}(\bm{x};h_{\hat{\theta}})=\min_{j\in[k],j\neq j_{1}}\|h_{\hat{\theta}}(\bm{x})-\bm{\mu}_{j}\|_{2}$ is the second smallest distance, and $j_{1}$ stands for the index of the closest centroid. By definition, we know $\ell_{\mathrm{KM}}(\bm{x})$ is always non-negative, and a smaller value of $\ell_{\mathrm{KM}}(\bm{x})$ indicates that the input $\bm{x}$ is more likely to be a boundary-adjacent point.
LCG-GMM. Similarly, we can define an informative guidance loss based on the coefficients of the Gaussian mixture models. Let $\{\bm{\mu}_{j},\mathbf{\Sigma}_{j})$ be the mean and covariance parameters returned by GMM with respect to the latent embeddings $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{l}\}$. Let $p_{j}(\bm{z})$ be the $j$-th GMM posterior probability defined by Equation 7. Then, the guidance loss is defined as:
[…]
Guided DDPM Fine-Tuning. To utilize our designed guidance losses in DDPM fine-tuning, we propose to first denoise $\bm{x}_{t}$ via a single DDPM reverse step $g_{\theta}(\bm{x}_{t},t)$, where $\bm{x}_{t}$ is a noisy image at $t$-th time step defined by the forward process, and then apply the respective guidance loss to $g_{\theta}(\bm{x}_{t},t)$. For any sample $\bm{x}_{0}\sim q_{0}$, time $t\sim\mathcal{U}(T)$ and noise $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, we define the following loss function $L_{\mathrm{reg}}(\theta)$ that will be used as an additional regularization term to fine-tune DDPM:
$L_{\mathrm{reg}}(\theta)=\mathbb{E}_{\bm{x}_{0},t,\bm{\epsilon}}\big[\ell_{\mathrm{guide}}\big(g_{\theta}(\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t)\big)\big],$ (15)
where $\ell_{\mathrm{guide}}$ corresponds to Equation 12 when we use prediction confidence, and refers to Equations 13 and 14 when using LCG-KM and LCG-GMM, respectively. Putting all the pieces together, the fine-tuning loss for direct generation of boundary-adjacent data is defined as:
$L_{\mathrm{tot}}(\theta)=L_{\mathrm{DDPM}}(\theta)+\lambda\cdot L_{\mathrm{reg}}(\theta),$ (16)
where $\lambda\geq 0$ balances the trade-off between high-quality image generation captured by the original DDPM loss and boundary-seeking data generation driven by our guidance loss. By penalizing the model weights through the regularization term, our method gradually pushes $\theta$ towards higher probabilities of generating boundary-adjacent data using stochastic gradient descent (SGD). Our later experiments show that when starting from a pre-trained DDPM (i.e., on CIFAR-10), a few fine-tuning epochs are sufficient to bias the diffusion model to generate the desirable subset of synthetic images.
IV Experiments
In this section, we test the performance of our data reduction strategies proposed in Section III compared with baseline SSAT algorithms. In particular, we first conduct experiments to evaluate our selection-based methods in terms of robustness and efficiency advantages when integrated into the SSAT pipeline (Section IV-A). Then, we highlight the computational advantages of our approaches, particularly their capabilities of accelerating the convergence and reducing the total runtime of SSAT (Section IV-B). Finally, we study the effectiveness of our DDPM fine-tuning schemes in further improving the efficiency of SSAT algorithms under a generative setup (Section IV-C). To begin with, we lay out the main experimental setup. All our experiments are conducted using a single A100 GPU. All the remaining details, including all the hyperparameters, for reproducing our experiments are provided in Appendix B.
Dataset. We test two image benchmarks in our evaluation, SVHN [26] and CIFAR-10 [27], both widely used in prior adversarial ML literature. We consider two scenarios in terms of the source of the utilized unlabeled dataset $\mathcal{S}_{u}$ for testing our selection-based strategies: (i) external–using an external dataset, and (ii) generative–using a synthetic dataset generated by some pre-trained generative model. For external scenarios, we follow the implementation protocols outlined by Carmon et al. [11]. To be more specific, SVHN models are trained on 7373K labeled digit images from the original dataset, supplemented by 531531K extra unlabeled SVHN images. For CIFAR-10, the original dataset comprises 5050K labeled training images and 1010K labeled test images, where models trained using SSAT are augmented by an external unlabeled dataset $\mathcal{S}_{u}$ with 500500K images sampled from the 8080M Tiny Images (80M-TI) dataset. For generative setup, we adopt 11M synthetic images generated by DDPM for both datasets, following the protocols documented by Gowal et al. [13].
Configuration. Following prior literature, we mainly consider $L_{\infty}$-norm bounded perturbations with $\epsilon=0.015$ on SVHN and $\epsilon=0.031$ on CIFAR-10. We adopt WideResNet (WRN) architectures and train models using TRADES [9] on both labeled and unlabeled data, coupled with pseudo labels returned by the intermediate model. In Appendix D, we evaluate the generalizability of our methods for varying $\ell_{\infty}$ perturbation size, vanilla PGD-based adversarial training with varying attack steps, and $\ell_{2}$ perturbations. To standardize training, an epoch is defined as processing 5050K data points, regardless of the total dataset size. Prior works also adopted this approach [11, 13], which ensures that training time per epoch remains consistent across datasets of varying sizes.
SSAT. For experiments with external unlabeled data, SSAT algorithms are trained for a total of $200$ epochs, whereas experiments with generated data extend to $400$ epochs. To compare the performance of our methods in a fair way, models are saved every $25$ epoch of SSAT, and the model achieving the highest robust accuracy is selected as the best model. This approach aligns with the early stopping practices commonly used in adversarial ML literature [9, 28]. In experiments with external data and ratio parameter $\alpha\in\{10\%,20\%\}$, we consistently observe peak robust accuracy around $100$ epochs for CIFAR-10 and $75$ epochs for SVHN. Conversely, when no data selection scheme is applied, peak performance is delayed to approximately $200$ epochs. Total training time is reported as the duration required to achieve the best model performance.
Intermediate Model. Our data selection and guided diffusion schemes require training an intermediate model $f_{\hat{\theta}}$, which we adopt the same architecture as the final model. This intermediate model is trained using standard supervised learning for $100$ epochs on a WRN-28-10 architecture, which takes around $55$ minutes. Notably, this time is excluded from the reported total runtime. The intermediate model fulfills two critical roles: it not only facilitates the pseudo-labeling of unlabeled data but also constitutes a necessary component for implementing any of our proposed strategies. As such, this training step is integral to the overall workflow, irrespective of whether a specific selection or fine-tuning method is employed.
Evaluation Metric. To test the robustness of the final model returned by SSAT, we mainly use multi-step PGD attacks (PGD) [7], due to their broad adoption in adversarial ML. We also report robust accuracies against AutoAttack (AA) [29] for more rigorous robustness evaluation and the model’s clean accuracy on normal inputs (Clean) as a reference.
TABLE I: SSAT performance under various unlabeled data selection schemes. We consider $\ell_{\infty}$ perturbations with $\epsilon=0.015$ for SVHN, and $\ell_{\infty}$ perturbations with $\epsilon=0.031$ for CIFAR-10. For each configuration, we underline the best robust accuracies.
@@ IV Experiments
LCS-KM $88.7$ $\underline{60.7}$ $\underline{57.8}$ $14.78$ $85.5$ $\underline{60.2}$ $\underline{57.2}$ $15.27$
IV-A Strategic Unlabeled Data Selection
@@
SVHN. Our evaluation results on SVHN are summarized in Table I, which demonstrate the efficacy of the proposed data selection schemes when employed in SSAT algorithms trained using a WRN-16-8 architecture with varying data ratios $\alpha\in\{1\%,10\%,20\%\}$, with comparisons to baseline SSAT performance with $\alpha\in\{0\%,100\%\}$. Specifically, employing our LCS-KM strategy with $10\%$ or $20\%$ of extra unlabeled data achieves robust accuracies of $86.3\%$ and $86.6\%$, respectively, which matches and even slightly outperforms the $86.0\%$ robustness achieved using the entire unlabeled dataset. In contrast, random selection with $20\%$ of the data yields a robust accuracy of only $82.3\%$, underscoring the superior efficacy of informed selection strategies such as LCS-KM. The robust accuracies evaluated using AutoAttack (i.e., the AA column of Table I) demonstrate similar patterns compared to those using multi-step PGD Attacks (i.e., the PGD column of Table I). Additionally, alternative methods, including PCS and LCS-GMM, consistently outperform random selection, confirming the effectiveness of our insights. Notably, the robustness of models trained using SSAT peaks when the boundary ratio parameter $\beta$ is inversely proportional to the selection ratio $\alpha$, suggesting that smaller $\beta$ values are favorable as the data inclusion proportion increases. We provide detailed sensitivity analyses of these hyperparameters in Section V-C.
@@
Our evaluations with 11M DDPM-generated data highlight that when using $10\%$ to $20\%$ of unlabeled data selected via LCS-KM, SSAT can achieve comparable robustness to the SSAT baseline. Since all SSAT algorithms listed under the column “DDPM-Generated” in Table I require the same amount of pre-generation time, we exclude it from the total runtime. In particular, training SSAT with reduced data requires only $75$ epochs to achieve the peak robust accuracy, while training with the full dataset requires up to $400$ epochs. Such a large reduction in computation time underscores the practical advantages of the proposed data selection strategies, which are further elaborated in Section IV-B. Besides, we observe in our experiments that using ground-truth labels reveals negligible differences compared to pseudo-labeled data, indicating that the primary performance gains stem from the selection strategy itself rather than label accuracy (see Table IV in Appendix C).
@@
CIFAR-10. We adopt a WRN-28-10 architecture for CIFAR-10 experiments. To assess the generalizability of our findings, we also conduct ablations on ResNet-18, which are detailed in Table V in Appendix C. Across varying ratios of utilized unlabeled data $\alpha$ and different data selection schemes, Table I reveals that selecting $20\%$ of the external data using LCS-KM attains robust accuracy of $60.7\%$, comparable to the $62.5\%$ obtained with the entire dataset. In comparison, random selection with the same ratio parameter achieves only $57.8\%$, underscoring the advantages of our strategic LCS technique with k-means. Similar to SVHN experiments, we extend our analysis to CIFAR-10 with synthetically generated data. Using 11M DDPM-generated CIFAR-10-like images, we discover that using just $20\%$ of the generated data via LCS-KM results in performance nearly identical to that achieved with the full dataset. These trends are also consistent across experiments with $\ell_{2}$ perturbations (see Figure 10 for supporting evidence).
@@
Our findings emphasize the robustness and adaptability of the proposed data selection schemes, especially for LCS-KM, even under varying experimental configurations. In terms of computational efficiency, CIFAR-10 experiments highlight the significant advantages of data selection. When leveraging selected external unlabeled data, optimal robust performance was achieved after $100$ epochs of training. By contrast, using the full unlabeled or generated datasets without selection required extended training durations of $200$ and $400$ epochs, respectively, to reach comparable performance. These results validate the efficacy of our latent clustering-based selection approach in enhancing semi-supervised adversarial training. The proposed method not only maintains the robust accuracy of the final model, $\theta_{\mathrm{final}}$, but also substantially reduces computational costs and accelerates the training process, underscoring its practical utility and scalability in SSAT frameworks.
(a) Vanilla AT
(b) Random Selection ($\alpha=10\%$)
(c) LCS-KM ($\alpha=10\%$)
(d) Full Utilization ($\alpha=100\%$)
Figure 1: Illustration of standard and robust accuracy curves of SSAT on CIFAR-10 with different configurations of unlabeled data selection from the external 500500K Tiny Images: (a) No extra data, (b) random selection with $\alpha=10\%$, (c) LCS-KM with $\alpha=10\%$, and (d) utilizing all 500500K unlabeled data.
TABLE II: Runtime comparison of SSAT with synthetically generated unlabeled data across various settings on CIFAR-10.
Ratio $\bm{\alpha}$ Method Fine-Tuning (h) Generation (h) Selection (h) #Epochs Total Time (h) PGD (%)
$0\%$ No Gen. $100$ $10.7$ $55.4$
$100\%$ Std. DDPM $3.90$ $400$ $61.0$ $61.8$
$20\%$ LCS-KM $3.90$ $0.97$ $100$ $19.1$ $60.3$
$20\%$ LCG-KM $0.43$ $0.77$ $100$ $15.7$ $60.2$
TABLE III: Performance of our DDPM fine-tuning strategies on CIFAR-10 under varying $\alpha$. The best robust accuracies are underlined.
Ratio $\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$) AA ($\%$)
$0\%$ No Gen. $84.9$ $55.4$ $49.2$
$100\%$ Std. DDPM $85.7$ $61.8$ $58.4$
$1\%$ PCG $84.4$ $55.4$ $52.8$
LCG-GMM $85.3$ $55.8$ $52.6$
LCG-KM $85.2$ $56.5$ $53.1$
$10\%$ PCG $84.7$ $56.7$ $54.2$
LCG-GMM $85.8$ $57.6$ $55.0$
LCG-KM $84.2$ $57.9$ $55.6$
$20\%$ PCG $85.9$ $58.3$ $55.9$
LCG-GMM $84.7$ $58.3$ $56.1$
LCG-KM $85.7$ $60.3$ $57.4$
IV-B Computational Benefits of Data Reduction